OpenStudy (anonymous):

What is the solution to the equation 14^ (x + 3) = 21 ? x ≈ –2.133 x ≈ –1.846 x ≈ 3.867 x ≈ 4.154

OpenStudy (cwrw238):

(x+3) ln 14 = ln 21 x+3 = ln 21 / ln 14 = 1.154 x = -2.154

OpenStudy (cwrw238):

oops - sorry its -4.154

OpenStudy (anonymous):

\[\Large \log(14)^{x+3}=\log(21)\] \[\Large (x+3)\cdot \log14=\log21\] \[\Large (x+3)=\frac{\log21}{log14}\] \[\Large x=\frac{\log21}{log14}-3\] use calculator and you'll get... \[\huge x\approx -1.84\]

OpenStudy (cwrw238):

what!! - i'm wrong the second time!! x = 1.154 - 3 = -1.846

OpenStudy (anonymous):

actually it is - 1.84. thank you.

OpenStudy (anonymous):

Here's proof

OpenStudy (cwrw238):

yw - i cant do simple algebra today!!! lol

OpenStudy (anonymous):

@cwrw238 It's important that you know to solve it (you know how it goes...) in calculations everyone makes mistakes ! .... that's normal. We're not robots!

OpenStudy (cwrw238):

right